Mannisto Inc. uses the FIFO inventory cost flow assumption. In a year of rising costs and prices, the firm reported net income of $1,500,000 and average assets of $10,000,000. If Mannisto had used the LIFO cost flow assumption in the same year, its cost of goods sold would have been $300,000 more than under FIFO , and its average assets would have been $300,000 less than under FIFO. A) Calculate the firm's ROI under each cost flow assumption (FIFO and LIFO). B) Suppose that two years later cost were $1,800,000 and $12,000,000, representatively. If LIFO had been used through the years, inventory values would have been $200,000 less than under FIFO. Calculate the firm's ROI under each cost flow assumption (FIFO and LIFO).
Answer #1

a) ROI =Net Income/Average Assets

FIFO= 1,500,000/10,000,000 = 15.0%
LIFO =1,200,000/9,700,000 = 12.4%

B)FIFO = 1,800,000/12,000,000 = 15%

LIFO = 1,600,000/11,800,000 =13.6%